a school administrator claims that less than 50% of the students of the school are dissatisfied by community cafeteria service. test this claim by using sample data obtained from a survey o 500 students of the school where 54% indicated their dissatisfaction of the community cafeteria service. use a =0.05


1
Expert's answer
2021-06-09T16:49:33-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0: p\geq0.5"

"H_1:p<0.5"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test is "z_c=-1.6449."

The rejection region for this left-tailed test is "R=\{z: z<-1.6449\}"


The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.54-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{500}}}\approx1.7889"

Since it is observed that "z=1.7889>-1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.5, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=P(z<1.7889)=0.9632," and since "p=0.9632>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.5, at the "\alpha=0.05" significance level.